### 04 Mar 2024 v1

@@ Abstract
Transformers have revolutionized computer vision and natural language processing, but their high computational complexity limits their application in high-resolution image processing and long-context analysis. This paper introduces Vision-RWKV (VRWKV), a model adap-ted from the RWKV model used in the NLP field with necessary modifications for vision tasks. Similar to the Vision Transformer (ViT), our model is designed to efficiently handle sparse inputs and demonstrate robust global processing capabilities, while also scaling up effectively, accommodating both large-scale parameters and extensive datasets. Its distinctive advantage lies in its reduced spatial aggregation complexity, which renders it exceptionally adept at processing high-resolution images seamlessly, eliminating the necessity for windowing operations. Our evaluations in image classification demonstrate that VRWKV matches ViT’s classification performance with significantly faster speeds and lower memory usage. In dense prediction tasks, it outperforms window-based models, maintaining comparable speeds. These results highlight VRWKV’s potential as a more efficient alternative for visual perception tasks. Code is released at https://github.com/OpenGVLab/Vision-RWKV.
@@ 1 Introduction
(3) Our model surpasses window-based ViTs and is comparable to global attention ViTs, demonstrating lower FLOPs and faster processing speeds as resolution increases. Notably, VRWKV-T achieves 75.1% top-1 accuracy trained only on the ImageNet-1K [8], outperforming DeiT-T [48] by 2.9 points. With large-scale parameters (i.e., 335M) and training data (i.e., ImageNet-22K), the top-1 accuracy of VRWKV-L is further boosted to 85.3%, which is slightly higher than ViT-L [13] (85.28 vs. missing 85.15). In addition, on COCO [29], a challenging downstream benchmark, our best model VRWKV-L achieves 50.6% box mAP, 1.9 points better than ViT-L (50.6 vs. missing 48.7).
@@ 2.2 Feature Aggregation Mechanism
As for the field of NLP, RNN-based operators [14, 36] have historically dominated because of their effectiveness in sequence modeling. RNNs and LSTMs excel in capturing temporal dependencies, making them suitable for tasks requiring an understanding of sequence dynamics. Subsequently, a significant shift occurred. The introduction of the transformer architecture [51] marked a turning point, with both NLP and computer vision fields shifting focus toward attention-based feature aggregation. The global attention mechanism overcomes the limitations of CNNs in modeling long-range dependencies and the shortcomings of RNNs in parallel computation while coming at a high computational cost.
@@
In addition to the efforts to optimize the global attention mechanism, various operators with linear complexity [37, 46, 16] have also been explored. For instance, RWKV [37] and RetNet [46] employed exponential decay to model global information efficiently while SSMs [17, 18, 42, 52] also exhibited linear complexity concerning sequence length and modification in Mamba [16] enable them to be input-dependent. Besides, XCA [1] achieved linear complexity by calculating the cross-variance between input tokens. However, the low efficiency of information interaction between tokens makes the need for the assistance of additional modules necessary to complete a comprehensive feature aggregation. Despite some concurrent efforts [31, 63, 15], adapting these NLP-derived techniques to vision tasks remains a challenge in maintaining stable training across larger and more complex vision models.
@@ 4.1 Image Classification
Settings. For -Tiny/Small/Base models, we conduct supervised training from scratch on ImageNet-1K [8]. Following the training strategy and data augmentation of DeiT [48], we use a batch size of 1024, AdamW [35] with a base learning rate of 5e-4, weight decay of 0.05, and cosine annealing schedule [34]. Images are cropped to the resolution of $224\times 224$ for training and validation. For the -Large models, we first pre-train them for 30 epochs on ImageNet-22K with a batch size of 4096 and resolution of $192\times 192$, and then fine-tune them for 20 epochs on ImageNet-1K to a higher resolution of $384\times 384$.
Results. We compare the results of our VRWKV with other hierarchical and non-hierarchical backbones on the ImageNet-1K dataset. As shown in Tab. 2, with the same number of parameters, computational complexity, and training/testing resolutions, VRWKV achieves comparable results to ViT. For example, compared to ViT-L, VRWKV-L achieves a similar top-1 accuracy of 85.3% at a resolution of $384\times 384$, with a slightly reduced computational cost. When the model size is smaller, VRWKV demonstrates higher baseline performance. In the case where both VRWKV-T and DeiT-T have FLOPs of 1.3G, VRWKV-T achieves a 2.9 point higher than DeiT-T. The exploration and utilization of linear attention mechanisms in VRWKV demonstrate its potential in vision tasks, making it a viable alternative to traditional ViT models that use global attention mechanisms. The comparable performance from tiny to large-size models also demonstrates that the VRWKV model possesses the scalability as ViT.
@@ 4.4 Ablation Study
Table 5: Ablation on key components of the proposed VRWKV. All models are trained from scratch on ImageNet-1K. The “original” means the original token shift in RWKV [37], which mixes the token in a single direction.
@@
MAE Pre-training. Similar to ViT, our VRWKV model can handle sparse inputs and benefits from the MAE pre-training [20]. By simply modifying the Q-Shift to perform a bidirectional shift operation, VRWKV can be pre-trained using MAE. The pre-trained weights can be directly fine-tuned for other tasks using a Q-Shift manner. Following the same MAE pre-training setting as ViT, and subsequent classification training similar to Sec. 4.1, our VRWKV-L achieves a top-1 accuracy improvement from 85.3% to 85.5% on ImageNet-1K val, showing the ability to acquire visual prior from masked image modeling.

### 07 Mar 2024 v2

@@ Abstract
Transformers have revolutionized computer vision and natural language processing, but their high computational complexity limits their application in high-resolution image processing and long-context analysis. This paper introduces Vision-RWKV (VRWKV), a model adap-ted from the RWKV model used in the NLP field with necessary modifications for vision tasks. Similar to the Vision Transformer (ViT), our model is designed to efficiently handle sparse inputs and demonstrate robust global processing capabilities, while also scaling up effectively, accommodating both large-scale parameters and extensive datasets. Its distinctive advantage lies in its reduced spatial aggregation complexity, which renders it exceptionally adept at processing high-resolution images seamlessly, eliminating the necessity for windowing operations. Our evaluations demonstrate that VRWKV surpasses ViT’s performance in image classification and has significantly faster speeds and lower memory usage processing high-resolution inputs. In dense prediction tasks, it outperforms window-based models, maintaining comparable speeds. These results highlight VRWKV’s potential as a more efficient alternative for visual perception tasks. Code is released at https://github.com/OpenGVLab/Vision-RWKV.
@@ 1 Introduction
(3) Our model surpasses window-based ViTs and is comparable to global attention ViTs, demonstrating lower FLOPs and faster processing speeds as resolution increases. Notably, VRWKV-T achieves 75.1% top-1 accuracy trained only on the ImageNet-1K [8], outperforming DeiT-T [49] by 2.9 points. With large-scale parameters (i.e., 335M) and training data (i.e., ImageNet-22K), the top-1 accuracy of VRWKV-L is further boosted to 86.0%, which is higher than ViT-L [12] (86.04 vs. missing 85.15). In addition, on COCO [27], a challenging downstream benchmark, our best model VRWKV-L achieves 50.6% box mAP, 1.9 points better than ViT-L (50.6 vs. missing 48.7).
@@ 2.2 Feature Aggregation Mechanism
As for the field of NLP, RNN-based operators [13, 34, 37] have historically dominated because of their effectiveness in sequence modeling. RNNs and LSTMs excel in capturing temporal dependencies, making them suitable for tasks requiring an understanding of sequence dynamics. Subsequently, a significant shift occurred. The introduction of the transformer architecture [52] marked a turning point, with both NLP and computer vision fields shifting focus toward attention-based feature aggregation. The global attention mechanism overcomes the limitations of CNNs in modeling long-range dependencies and the shortcomings of RNNs in parallel computation while coming at a high computational cost.
@@
In addition to the efforts to optimize the global attention mechanism, various operators with linear complexity [35, 47, 15, 36, 46] have also been explored. For instance, RWKV [35] and RetNet [47] employed exponential decay to model global information efficiently while SSMs [16, 17, 42, 53] also exhibited linear complexity concerning sequence length and modification in Mamba [15] enable them to be input-dependent. Besides, XCA [1] achieved linear complexity by calculating the cross-variance between input tokens. However, the low efficiency of information interaction between tokens makes the need for the assistance of additional modules necessary to complete a comprehensive feature aggregation. Despite some concurrent efforts [29, 63, 14], adapting these NLP-derived techniques to vision tasks remains a challenge in maintaining stable training across larger and more complex vision models.
@@ 4.1 Image Classification
Settings. For -Tiny/Small/Base models, we conduct supervised training from scratch on ImageNet-1K [8]. Following the training strategy and data augmentation of DeiT [49], we use a batch size of 1024, AdamW [33] with a base learning rate of 5e-4, weight decay of 0.05, and cosine annealing schedule [32]. Images are cropped to the resolution of $224\times 224$ for training and validation. For the -Large models, we first pre-train them for 90 epochs on ImageNet-22K with a batch size of 4096 and resolution of $192\times 192$, and then fine-tune them for 20 epochs on ImageNet-1K to a higher resolution of $384\times 384$.
Results. We compare the results of our VRWKV with other hierarchical and non-hierarchical backbones on the ImageNet-1K validation dataset. As shown in Tab. 2, with the same number of parameters, computational complexity, and training/testing resolutions, VRWKV achieves better results than ViT. For example, when VRWKV-T has slightly lower FLOPs than DeiT-T(1.2G vs. 1.3G), our VRWKV-T achieves a 2.9 point higher than DeiT-T on top-1 accuracy. When the model size scales up, VRWKV still demonstrates higher baseline performance. In the case of large models, VRWKV-L achieves a 0.8 point higher top-1 accuracy of 86.0% at the resolution of $384\times 384$ than ViT-L, with a slightly reduced computational cost. The superior performance from tiny to large-size models demonstrates that the VRWKV model possesses the scalability as ViT. Additionally, after using Bamboo-47K [61] in the pre-train process, the performance of VRWKV-L can be further boosted to 86.5%, indicating that our VRWKV also possesses the ability like ViT to benefit from pre-training on large-scale datasets. The exploration of VRWKV in classification tasks demonstrates its potential to be a viable alternative to traditional ViT models.
@@ 4.4 Ablation Study
Table 5: Ablation on key components of the proposed VRWKV. All models are trained from scratch on ImageNet-1K. The “original” means the original token shift in RWKV [35] which mixes the token in a single direction is used.
@@
MAE Pre-training. Similar to ViT, our VRWKV model can handle sparse inputs and benefits from the MAE pre-training [19]. By simply modifying the Q-Shift to perform a bidirectional shift operation, VRWKV can be pre-trained using MAE. The pre-trained weights can be directly fine-tuned for other tasks using a Q-Shift manner. Following the same MAE pre-training setting as ViT, and subsequent classification training similar to Sec. 4.1, our VRWKV-L achieves a top-1 accuracy improvement from 86.0% to 86.2% on ImageNet-1K val, showing the ability to acquire visual prior from masked image modeling.